OpenStudy (anonymous):

Which represent the first two terms of the sequence: a1 = 2 and an = 4(an -1)2 ?

OpenStudy (anonymous):

2, 16?

Nnesha (nnesha):

n = next term term that you have to find

Nnesha (nnesha):

a_1 = first term so yes first term is 2 2nd term ?

OpenStudy (anonymous):

64; 65,536

Nnesha (nnesha):

nope first one is right

Nnesha (nnesha):

that suppose to be 4(n -1)2

OpenStudy (xapproachesinfinity):

it is mysterious how you wrote it! is that \[\rm a_n=4(n-1)^2\]

OpenStudy (xapproachesinfinity):

seems you already got the first term ( don't even why they asked you do it again) the second term plug in n=2

OpenStudy (anonymous):

-16, 1024 so wouldn't this be the correct answer

OpenStudy (xapproachesinfinity):

first tell us how that sequence written it is still mysterious to us

OpenStudy (xapproachesinfinity):

taking a picture is a good alternative if you don't know how to write it here

OpenStudy (xapproachesinfinity):

do you know what i'm talking about?

OpenStudy (xapproachesinfinity):

i mean, please rewrite this "an = 4(an -1)2" accurately it is not well written and i don't know what that is it is vague and we cannot do anything if we don't know what that is

OpenStudy (xapproachesinfinity):

if you have the problem in your computer screenshot it and post if not take a picture of by phone and post it

OpenStudy (anonymous):

a1 = 2 and an = 4(an 1)2 ?

OpenStudy (xapproachesinfinity):

i'm saying that does not make any sense write it the way it is written in your paper, \[\rm a_n=4(a_n-1)2\] this is not good understand!! that's what you wrote

OpenStudy (xapproachesinfinity):

i wish i have some magic power to see what was the question written like

OpenStudy (anonymous):

That's what the question is sir!

OpenStudy (xapproachesinfinity):

i don't think so, i'm sure it is something else i'm saying through some experience. the way you wrote is invalid and has no meaning what so ever!

OpenStudy (anonymous):

What represents the FIRST TERMS of the sequence

OpenStudy (xapproachesinfinity):

i'm sure you are not writing the equation correctly!!

OpenStudy (anonymous):

Now does it make sense?

OpenStudy (anonymous):

Good gravy! im getting stressed out!

OpenStudy (xapproachesinfinity):

you still didnt write the equation do you understand what i'm talking im talking about an=4(an-1)2 this one stir at it and write it exactly like they wrote it

OpenStudy (xapproachesinfinity):

there is no reason to get stressed, relax :)

OpenStudy (anonymous):

Ok, lets try another one instead...

OpenStudy (anonymous):

Which graph represents the sequence: 2, 3.5, 5, 6.5, 8, ...?

OpenStudy (xapproachesinfinity):

and where are the graphs?

OpenStudy (anonymous):